# Folded multistability and hidden critical point in microwave-driven Rydberg atoms

### v2

@@
11footnotetext: Y.M, B.L, and L.H.Z contribute equally to this work.22footnotetext: dds@ustc.edu.cn
Folded multistability and hidden critical point in microwave-driven Rydberg atoms
Yu Ma1,2,⋆    Bang Liu1,2,⋆    Li-Hua Zhang1,2,⋆    Ya-Jun Wang1,2    Zheng-Yuan Zhang1,2    Shi-Yao Shao1,2    Qing Li1,2    Han-Chao Chen1,2    Jun Zhang1,2    Tian-Yu Han1,2    Qi-Feng Wang1,2    Jia-Dou Nan1,2    Yi-Ming Yin1,2    Dong-Yang Zhu1,2    Bao-Sen Shi1,2    Dong-Sheng Ding1,2,† 1Key Laboratory of Quantum Information, University of Science and Technology of China, Hefei, Anhui 230026, China. 2Synergetic Innovation Center of Quantum Information and Quantum Physics, University of Science and Technology of China, Hefei, Anhui 230026, China.
(September 9, 2024)
@@
Due to exaggerated properties of Rydberg atoms [1, 2, 3, 4], the Rydberg atoms systems have been shown to exhibit many intriguing characteristics of complex systems, including non-equilibrium phase transitions [5, 6, 7, 8], self-organization [9, 10, 11], non-ergodicity and time crystals [12, 13, 14, 15, 16, 17, 18], and non-Hermitian many-body physics [19]. Because of interaction-induced nonlinearity, the system settles into one of two stable configurations, depending on the external influences such as electric fields or the density of Rydberg atoms [20, 21, 18, 22], resulting in complex bistability [6, 10] or hysteresis loop [19]. The stationary states within the bistable regime can be effectively described using a double-well potential [23]. This framework encapsulates the transitions between a high-density phase and a low-density phase. The emergence of bistability in Rydberg atoms has become an intriguing resource to study non-equilibrium many-body physics [6, 7, 24, 20, 25]. In contrast to other systems there is no optical cavity feedback [26, 27, 28]; or requirement for cryogenic temperatures [29].
It is worth noting that the multiple phases could fold together when Rydberg atoms are driven by microwave fields, thus the system cannot be broken down to just two global states. Within a hysteresis loop, there can be multiple states folded on top of each other, resulting from high-order nonlinearity and high-order symmetry breaking. This phenomenon is often observed in complex systems characterized by non-linear dynamics, feedback loops, and multiple interconnected variables. As parameters are varied within a certain range, different states emerge and coexist, resulting in complex hysteresis loops and multistability across various variables in different systems, including biological systems, chemical reactions, ecosystems and climate dynamics [30, 31, 32, 33, 34]. This multistable effect has been predicted and investigated in coupled atom-cavity systems [35, 36, 37, 38] and a semiconductor microcavity [39, 40, 41, 42]. The mutually fold in a multistable system could exhibit bistability within the hysteresis loop, which results in hidden multistability where the number of output states is less than that of the steady states of system. This makes it challenging to identify all possible states in systems that contain hidden multistability, making it a compelling topic for exploration and theoretical investigation [43].
@@
In this work, we observe a phase transition from bistability to multistability in a thermal Cesium Rydberg atoms system under a microwave field driving. At low microwave field intensities, the strong interaction between Rydberg atoms induces bistability. However, when the microwave field intensity is increased, the Rydberg energy levels undergo splitting due to the Autler-Townes effect [44, 45]. This breaks $\mathbb{Z}_{3}$-symmetry and results in a phase transition from bistability to multistability. This process reveals a hidden critical point, near which a third, folded state is concealed within the bistability hysteresis loop (which only presents two observable states). Consequently, multiple non-equilibrium steady states coexist within the same parameter range. In this context, the system’s behavior cannot be reduced to merely two stable states. By carefully tuning the system parameters, we demonstrate that these multiple non-equilibrium steady states can be distinctly identified and distinguished, including hidden states and sublevels-dependent steady states. The reported findings on the phase transition and the folding effect of multistability in the Rydberg atoms system hold significant promise for studying and characterizing intriguing non-equilibrium phases in the realms of nonlinearity and many-body physics.
Physical model
@@
To demonstrate the phase transition from the state in bistability to that in multistability, we build a simplified physical model to uncover the underlying mechanism of experimental observations. The model consists of $N$ interacting three-level atoms with a ground state |$\ket{g}$ and the Rydberg states |$\ket{r_{1}}$ and |$\ket{r_{2}}$ (with a decay rate of $\gamma_{1}$ and $\gamma_{2}$). The atoms are coupled by a laser with a Rabi frequency $\Omega$ and detuning $\Delta_{1}$. For the experiment, this process can be realized by an off-resonance two-photon Raman excitation scheme, see more details in Methods section. With a large detuning $\Delta_{p}$, the intermediate state |$\ket{e}$ is adiabatically eliminated. The Rydberg states |$\ket{r_{1}}$ and |$\ket{r_{2}}$ are driven by a resonant microwave field with a Rabi frequency $\Omega_{m}$, a frequency $\omega_{m}$, and a detuning $\Delta_{2}$.
@@
The energy diagram of the model is depicted in Fig.1(a). The experimental setup can be found in Fig.1(b), in which the insert figure is the measured transmission spectrum with two jumps. The two jumps exhibit the characteristic of multistability. To describe the underlying mechanism behind the observations, we plot different potential wells to explain it. Here, one jump corresponds to the transition between two stationary states in a double-well potential [23], while two jumps reveal the transitions in a triple-well potential, as illustrated in Fig.1(c). The interaction between Rydberg atoms is governed by the van der Waals interaction strength $V={C_{6}}/{r^{6}}$ [where $C_{6}$ is coefficient and $r$ represents the distance between the Rydberg atoms]. The Hamiltonian of system is:
@@
We calculate the steady-state solution of the Lindblad master equation $\dot{\rho}=-i[\hat{H},\rho]+\mathcal{L}[\rho]$ under the condition $\dot{\rho}=0$, employing the mean-field approximation $\Delta_{1}\rightarrow\Delta_{1}+\Delta_{\text{shift}}-V\rho_{r_{1}r_{1}}$, see more information in Methods section. Here, $\mathcal{L}[\rho]$ is the Lindblad term, $\Delta_{\text{shift}}$ represents the microwave fields-dependent energy shift, which is proportional to the Rabi frequency of the microwave field. For simplicity, we set $\Delta_{\text{shift}}=\lambda\Omega_{m}$ with $\lambda=1.8$. Consequently, we obtain a series of real roots for the Rydberg population $\rho_{r_{1}r_{1}}$ as a function of $\Delta_{1}$.
Figure 2: Theoretical phase diagram and multistability. (a) Theoretical phase diagram depicting the variation of $\rho_{r_{1}r_{1}}$ with respect to microwave field power and laser detuning under forward and backward scanning detuning. The colored boundary lines on the left side of the figure correspond to the boundaries of the phase diagram. As the microwave field power gradually increases, the population difference $\rho_{r_{1}r_{1}}$ in the bistability regime becomes split, eventually leading to the emergence of multistability. The color bar is to show the mapping of values to colors in the displayed data. (b) The spectrum of $\rho_{r_{1}r_{1}}$ is presented for varying microwave field intensities. The blue and red lines show the forward and backward scanning detuning, respectively. The lower panel of (b) illustrates the bistability range, while the middle and upper panels showcase the spectrum in a multistable regime. The dashed lines in these figures indicate the calculated solutions obtained from the Lindblad master equation. In these calculations, the parameters are set as $\Gamma_{1}=8\gamma$, $\Gamma_{2}=\gamma$, $\Delta_{2}=1.5\gamma$, and $V=-200\gamma$.
Phase diagram and multistability
@@
The physical picture of system is captured by a mean-field potential in dissipative Rydberg atoms [23]. When the Rydberg atoms interact with weak microwave fields, the $\mathbb{Z}_{2}$-symmetry is broken because of strong interaction between Rydberg atoms, leading to a nonequilibrium phase transition between two distinct stationary states of system in a self-organization configuration [6, 10]. This corresponds to a single jump in the spectrum associated with a transition between each well in the double-well potential, as shown in the upper panel of Fig. 1(c). The multistability indicates the presence of a higher-order phase transition, which involves the breaking of $\mathbb{Z}_{3}$-symmetry, allowing the system to settle into one of three distinct states. When scanning $\Delta_{1}$, we observe two jumps in the spectrum, corresponding to two transitions within a triple-well potential, as depicted in the lower panel of Fig. 1(c).
Figure 3: Measured phase diagram and multistability. (a) Measured phase diagram of difference of probe transmission versus the microwave field power and laser detuning under forward and backward scanning detuning. Here ”Microwave Power” refers to the output microwave power on settings of signal generator. As the microwave power increases, the transmission difference under two scanning directions becomes split gradually from a sharp band. The color bar is to show the mapping of values to colors in the displayed data. (b) The spectrum of probe light with different microwave field power. Blue and red lines correspond to forward and backward scanning detuning. Spectra in (b) are extracted as representation of different transition stages from (a) at different microwave powers. In this process, we set the frequency of the microwave field as $\omega_{m}=2\pi\times$ 26.17 GHz.
In the experiment, we utilized a two-photon electromagnetically induced transparency (EIT) scheme [46, 47, 48] to excite atoms to a Rydberg state in a cesium thermal vapor, as shown in the experimental configurations in Fig. 1(b). Further information regarding the laser beams and parameters of microwave fields can be found in the Methods section. The population of Rydberg atoms can be monitored by the response of the atoms to the probe field, allowing for the measurement of the EIT spectrum under various conditions. Bistability and multistability can be achieved by scanning the detuning of coupling field in both forward and backward directions. The regimes of bistability and multistability are determined by subtracting the spectrum from the backward scan and that from the forward scan $\delta T=T^{B}-T^{F}$, where $T^{B}$ and $T^{F}$ are the corresponding probe transmission under backward and forward scanning $\Delta_{1}$.
@@
Figure 4: Hidden critical point and multistability. (a) Calculated EIT spectrum under forward (blue) and backward (red) scanning $\Delta_{1}$. The population lines under these two scan cases form a bistable hysteresis loop. Within the bistability regime, there is a hidden multistable state that cannot be obtained by scanning $\Delta_{1}$. (b) The phase diagram obtained by combining $\delta\Delta_{1}$ with $\rho_{r_{1}r_{1}}$ having $\geq 3$ roots (blue circle) and $\geq 4$ roots (red circle) respectively, here, $\delta\Delta_{1}$ represents the maximal detuning difference for multiple roots. The region of pure bistability (with $\mathbb{Z}_{2}$-symmetry breaking) is indicated by the gray area, while the region of multistability (with $\mathbb{Z}_{3}$-symmetry breaking) is highlighted in green and red. The concealed multistable state appears as bistability in observations denoted by the green color. (c) Measured hidden multistable state. The hidden state [marked by the shaded arrow] can be identified by scanning the detuning at the starting point $\Delta_{1}$. The trajectory of probe transmission is shown by the green color, in which there is a hidden jump that corresponds to the phase transition from the hidden state to the unhidden state (with a high transmission level). (d) Measured the phase transitions with varying the microwave power $P=1$ dBm to $P=-17$ dBm. The left panel is the phase diagram, and the right panel represents the transmissions under the microwave power $P=-1$ dBm, -5 dBm, -9 dBm, and -14 dBm, respectively. The point where the phase transition disappears corresponds to the hidden critical point. The color bar is to show the mapping of values to colors in the displayed data. Based on these results, we can identify a hidden critical point at which the $\mathbb{Z}_{3}$-symmetry begins to break. In this process, we set the frequency of the microwave field as $\omega_{m}=2\pi\times$ 26.12 GHz.
Hidden critical point
@@
Multiple jumps in spectrum
@@
All phase transition points in the phase diagram exhibit oscillatory effects versus the microwave frequency $\omega_{m}$. Figure 5(c) illustrates the dynamics of the first phase transition point, showcasing a pronounced oscillation effect. The variations in $\omega_{m}$ induce resonance and off-resonance conditions between the microwave field and the Rydberg sublevels of $m_{j}$, resulting in distinct population distributions. This leads to jumps at different detunings, culminating in the observed oscillatory phenomenon. In addition, we present a transmission spectrum plotted as a function of $\omega_{m}$ at a detuning of $\Delta_{1}=-2\pi\times 750$ MHz, as shown in Fig 5(d). Here, several counterintuitive phase transition points can be observed, where the transmission exhibits abrupt increases and decreases, resulting in all-microwave multistate switching [38]. This behavior arises from the varying coupling strengths under both resonance and off-resonance conditions, as discussed above.
@@
Discussions
Studying microwave-driven multistability and their transitions in Rydberg atom systems offers significant insights in other aspects induced bistability or multistability both theoretically and experimentally [5, 23, 49, 24, 50, 51]. For instance, random particle motion may also drive non-equilibrium phase transitions [24]. In addition, quantum fluctuations can facilitate transitions between multiple stable collective states, as suggested by classical mean-field theory [5] and the variational principle [49]. In these methods and technologies, the effective coupling between Rydberg atoms and microwave fields enables more precise control over phase transitions and multistability.
Furthermore, exploring hidden critical points in Rydberg atom systems could illuminate broader questions within driven-dissipative many-body systems, particularly in the context of high-order symmetry breaking, including applications of identifying some exotic hidden phases [43]. As a quantum many-body system may have multiple possible hidden critical points, it is promising to study the mechanism of these hidden critical points [52]. The developed methods to find the hidden critical points in this work allow us to uncover and understand previously unknown phases of matter and unlocking new frontiers in quantum many-body physics.

### v3

@@
Abstract
@@
I Introduction
Due to exaggerated properties of Rydberg atoms Saffman et al. (2010); Firstenberg et al. (2016); Adams et al. (2019); Browaeys and Lahaye (2020), the Rydberg atoms systems have been shown to exhibit many intriguing characteristics of complex systems, including non-equilibrium phase transitions Lee et al. (2012); Carr et al. (2013); Malossi et al. (2014); Lesanovsky and Garrahan (2014), self-organization Helmrich et al. (2020); Ding et al. (2020); Klocke et al. (2020), non-ergodicity and time crystals Gambetta et al. (2019); Ding et al. (2024); Wadenpfuhl and Adams (2023); Wu et al. (2024); Liu et al. (2024a, b, c), and non-Hermitian many-body physics Zhang et al. (2024a). Because of interaction-induced nonlinearity, the system settles into one of two stable configurations, depending on the external influences such as electric fields or the density of Rydberg atoms (Wade et al., 2018; Ding et al., 2022; Liu et al., 2024c; Zhang et al., 2024b), resulting in complex bistability Carr et al. (2013); Ding et al. (2020) or hysteresis loop Zhang et al. (2024a). The stationary states within the bistable regime can be effectively described using a double-well potential (Marcuzzi et al., 2014). This framework encapsulates the transitions between a high-density phase and a low-density phase. The emergence of bistability in Rydberg atoms has become an intriguing resource to study non-equilibrium many-body physics (Carr et al., 2013; Malossi et al., 2014; Šibalić et al., 2016; Wade et al., 2018; Weller et al., 2019). In contrast to other systems there is no optical cavity feedback (Gibbs et al., 1976; Wang et al., 2001a, b); or requirement for cryogenic temperatures (Hehlen et al., 1994).
It is worth noting that the multiple phases could fold together when Rydberg atoms are driven by microwave fields, thus the system cannot be broken down to just two global states. Within a hysteresis loop, there can be multiple states folded on top of each other, resulting from high-order nonlinearity and high-order symmetry breaking. This phenomenon is often observed in complex systems characterized by non-linear dynamics, feedback loops, and multiple interconnected variables. As parameters are varied within a certain range, different states emerge and coexist, resulting in complex hysteresis loops and multistability across various variables in different systems, including biological systems, chemical reactions, ecosystems and climate dynamics May (1977); Simoyi et al. (1982); Paillard (1998); Laurent and Kellershohn (1999); Pisarchik and Feudel (2014). This multistable effect has been predicted and investigated in coupled atom-cavity systems (Kitano et al., 1981; Cecchi et al., 1982; Joshi and Xiao, 2003; Sheng et al., 2012) and a semiconductor microcavity (Gippius et al., 2007; Paraïso et al., 2010; Cerna et al., 2013; Goblot et al., 2019). The mutually fold in a multistable system could exhibit bistability within the hysteresis loop, which results in hidden multistability where the number of output states is less than that of the steady states of system. This makes it challenging to identify all possible states in systems that contain hidden multistability, making it a compelling topic for exploration and theoretical investigation Bi et al. (2024).
@@ I Introduction
In this work, we observe a phase transition from bistability to multistability in a thermal Cesium Rydberg atoms system under a microwave field driving. At low microwave field intensities, the strong interaction between Rydberg atoms induces bistability. However, when the microwave field intensity is increased, the Rydberg energy levels undergo splitting due to the Autler-Townes effect Sedlacek et al. (2012); Liu et al. (2023). This breaks $\mathbb{Z}_{3}$-symmetry and results in a phase transition from bistability to multistability. This process reveals a hidden critical point, near which a third, folded state is concealed within the bistability hysteresis loop (which only presents two observable states). Consequently, multiple non-equilibrium steady states coexist within the same parameter range. In this context, the system’s behavior cannot be reduced to merely two stable states. By carefully tuning the system parameters, we demonstrate that these multiple non-equilibrium steady states can be distinctly identified and distinguished, including hidden states and sublevels-dependent steady states. The reported findings on the phase transition and the folding effect of multistability in the Rydberg atoms system hold significant promise for studying and characterizing intriguing non-equilibrium phases in the realms of nonlinearity and many-body physics.
II Physical model
@@
To demonstrate the phase transition from the state in bistability to that in multistability, we build a simplified physical model to uncover the underlying mechanism of experimental observations. The model consists of $N$ interacting three-level atoms with a ground state |$\ket{g}$ and the Rydberg states |$\ket{r_{1}}$ and |$\ket{r_{2}}$ (with a decay rate of $\gamma_{1}$ and $\gamma_{2}$). The atoms are coupled by a laser with a Rabi frequency $\Omega$ and detuning $\Delta_{1}$. For the experiment, this process can be realized by an off-resonance two-photon Raman excitation scheme, see more details in appendix. With a large detuning $\Delta_{p}$, the intermediate state |$\ket{e}$ is adiabatically eliminated. The Rydberg states |$\ket{r_{1}}$ and |$\ket{r_{2}}$ are driven by a resonant microwave field with a Rabi frequency $\Omega_{m}$, a frequency $\omega_{m}$, and a detuning $\Delta_{2}$.
@@
The energy diagram of the model is depicted in Fig.1(a). The experimental setup can be found in Fig.1(b), in which the insert figure is the measured transmission spectrum with two jumps. The two jumps exhibit the characteristic of multistability. To describe the underlying mechanism behind the observations, we plot different potential wells to explain it. Here, one jump corresponds to the transition between two stationary states in a double-well potential (Marcuzzi et al., 2014), while two jumps reveal the transitions in a triple-well potential, as illustrated in Fig.1(c). The interaction between Rydberg atoms is governed by the van der Waals interaction strength $V={C_{6}}/{r^{6}}$ [where $C_{6}$ is coefficient and $r$ represents the distance between the Rydberg atoms]. The Hamiltonian of system is:
@@
We calculate the steady-state solution of the Lindblad master equation $\dot{\rho}=-i[\hat{H},\rho]+\mathcal{L}[\rho]$ under the condition $\dot{\rho}=0$, employing the mean-field approximation $\Delta_{1}\rightarrow\Delta_{1}+\Delta_{\text{shift}}-V\rho_{r_{1}r_{1}}$, see more information in appendix. Here, $\mathcal{L}[\rho]$ is the Lindblad term, $\Delta_{\text{shift}}$ represents the microwave fields-dependent energy shift, which is proportional to the Rabi frequency of the microwave field. For simplicity, we set $\Delta_{\text{shift}}=\lambda\Omega_{m}$ with $\lambda=1.8$ as an ionization-induced shift coefficient. Consequently, we obtain a series of real roots for the Rydberg population $\rho_{r_{1}r_{1}}$ as a function of $\Delta_{1}$.
Figure 2: Theoretical phase diagram and multistability. (a) Theoretical phase diagram depicting the variation of $\rho_{r_{1}r_{1}}$ with respect to microwave field power and laser detuning under forward and backward scanning detuning. The colored boundary lines on the left side of the figure correspond to the boundaries of the phase diagram. As the microwave field power gradually increases, the population difference $\rho_{r_{1}r_{1}}$ in the bistability regime becomes split, eventually leading to the emergence of multistability. The color bar is to show the mapping of values to colors in the displayed data. (b) The spectrum of $\rho_{r_{1}r_{1}}$ is presented for varying microwave field intensities with dotted lines. The blue and red lines show the behaviors of forward and backward scanning detuning, respectively, while the dotted lines that are not covered by solid lines at all represents solutions to unstable states. The lower panel of (b) illustrates the bistability range, while the middle and upper panels showcase the spectrum in a multistable regime. The dashed lines in these figures indicate the calculated solutions obtained from the Lindblad master equation. In these calculations, the parameters are set as $\Gamma_{1}=8\gamma$, $\Gamma_{2}=\gamma$, $\Delta_{2}=1.5\gamma$, and $V=-200\gamma$.
III Phase diagram and multistability
@@
The physical picture of system is captured by a mean-field potential in dissipative Rydberg atoms (Marcuzzi et al., 2014). When the Rydberg atoms interact with weak microwave fields, the $\mathbb{Z}_{2}$-symmetry is broken because of strong interaction between Rydberg atoms, leading to a nonequilibrium phase transition between two distinct stationary states of system in a self-organization configuration Carr et al. (2013); Ding et al. (2020). This corresponds to a single jump in the spectrum associated with a transition between each well in the double-well potential, as shown in the upper panel of Fig. 1(c). The multistability indicates the presence of a higher-order phase transition, which involves the breaking of $\mathbb{Z}_{3}$-symmetry, allowing the system to settle into one of three distinct states. When scanning $\Delta_{1}$, we observe two jumps in the spectrum, corresponding to two transitions within a triple-well potential, as depicted in the lower panel of Fig. 1(c).
Figure 3: Measured phase diagram and multistability. (a) Measured phase diagram of difference of probe transmission versus the microwave field power and laser detuning under forward and backward scanning detuning. Here ”Microwave Power” refers to the output microwave power on the signal generator settings . As the microwave power increases, the transmission difference under two scanning directions becomes split gradually from a sharp band. The color bar is to show the mapping of values to colors in the displayed data. (b) The spectrum of probe light with different microwave field power. Blue and red lines correspond to forward and backward scanning detuning. Spectra in (b) are extracted as representation of different transition stages from (a) at different microwave powers. In this process, we set the frequency of the microwave field as $\omega_{m}=2\pi\times$ 26.17 GHz.
In the experiment, we utilized a two-photon electromagnetically induced transparency (EIT) scheme Harris et al. (1990); Mohapatra et al. (2007); Petrosyan et al. (2011) to excite atoms to a Rydberg state in a cesium thermal vapor, as shown in the experimental configurations in Fig. 1(b). Further information regarding the laser beams and parameters of microwave fields can be found in appendix. The population of Rydberg atoms can be monitored by the response of the atoms to the probe field, allowing for the measurement of the EIT spectrum under various conditions. Bistability and multistability can be achieved by scanning the detuning of coupling field in both forward and backward directions. The regimes of bistability and multistability are determined by subtracting the spectrum from the backward scan and that from the forward scan $\delta T=T^{B}-T^{F}$, where $T^{B}$ and $T^{F}$ are the corresponding probe transmission under backward and forward scanning $\Delta_{1}$.
@@
Figure 4: Hidden critical point and multistability. (a) Calculated EIT spectrum under forward (blue) and backward (red) scanning $\Delta_{1}$. The population lines under these two scan cases form a bistable hysteresis loop. Within the bistability regime, there is a hidden multistable state that cannot be obtained by scanning $\Delta_{1}$. (b) Scatter plot of $\delta\Delta_{1}$ with $\rho_{r_{1}r_{1}}$ having $\geq 3$ roots (blue circle) and $\geq 4$ roots (red circle) respectively against microwave power, and here $\delta\Delta_{1}$ represents the whole span of $\Delta_{1}$ with certain number of roots. The region of pure bistability (with $\mathbb{Z}_{2}$-symmetry breaking) is indicated by the gray area, while the region of multistability (with $\mathbb{Z}_{3}$-symmetry breaking) is highlighted in green and red. The concealed multistable state appears as bistability in observations denoted by the green color. (c) Measured hidden multistable state. The hidden state [marked by the shaded arrow] can be identified by scanning the detuning at the starting point $\Delta_{1}$. The trajectory of probe transmission is shown by the green color, in which there is a hidden jump that corresponds to the phase transition from the hidden state to the unhidden state (with a high transmission level). (d) Measured the phase transitions with varying the microwave power $P=1$ dBm to $P=-17$ dBm. The left panel is the phase diagram, and the right panel represents the transmissions under the microwave power $P=-1$ dBm, -5 dBm, -9 dBm, and -14 dBm, respectively. The point where the phase transition disappears corresponds to the hidden critical point. The color bar is to show the mapping of values to colors in the displayed data. Based on these results, we can identify a hidden critical point at which the $\mathbb{Z}_{3}$-symmetry begins to break. In this process, we set the frequency of the microwave field as $\omega_{m}=2\pi\times$ 26.12 GHz.
IV Hidden critical point
@@
V Multiple jumps in spectrum
@@
All phase transition points in the phase diagram exhibit oscillatory effects versus the microwave frequency $\omega_{m}$. Figure 5(c) illustrates the dynamics of the first phase transition point, showcasing a pronounced oscillation effect. The variations in $\omega_{m}$ induce resonance and off-resonance conditions between the microwave field and the Rydberg sublevels of $m_{j}$, resulting in distinct population distributions. This leads to jumps at different detunings, culminating in the observed oscillatory phenomenon. In addition, we present a transmission spectrum plotted as a function of $\omega_{m}$ at a detuning of $\Delta_{1}=-2\pi\times 750$ MHz, as shown in Fig 5(d). Here, several counterintuitive phase transition points can be observed, where the transmission exhibits abrupt increases and decreases, resulting in all-microwave multistate switching Sheng et al. (2012). This behavior arises from the varying coupling strengths under both resonance and off-resonance conditions, as discussed above.
@@
VI Discussions
Studying microwave-driven multistability and their transitions in Rydberg atom systems offers significant insights in other aspects induced bistability or multistability both theoretically and experimentally (Lee et al., 2012; Marcuzzi et al., 2014; Weimer, 2015; Šibalić et al., 2016; Levi et al., 2016; Ding et al., 2021). For instance, random particle motion may also drive non-equilibrium phase transitions (Šibalić et al., 2016). In addition, quantum fluctuations can facilitate transitions between multiple stable collective states, as suggested by classical mean-field theory (Lee et al., 2012) and the variational principle (Weimer, 2015). In these methods and technologies, the effective coupling between Rydberg atoms and microwave fields enables more precise control over phase transitions and multistability.
Furthermore, exploring hidden critical points in Rydberg atom systems could illuminate broader questions within driven-dissipative many-body systems, particularly in the context of high-order symmetry breaking, including applications of identifying some exotic hidden phases Bi et al. (2024). As a quantum many-body system may have multiple possible hidden critical points, it is promising to study the mechanism of these hidden critical points Haldar et al. (2023). The developed methods to find the hidden critical points in this work allow us to uncover and understand previously unknown phases of matter and unlocking new frontiers in quantum many-body physics.